This is from an on-line quiz, so unfortunately I do not have an ISBN number for you.
Scott runs 3 miles in 26 minutes. At the same rate, how many miles can he run in 39 minutes?
I divided 3 into 26 to get 8.66, but then I don't know???? I would appreciate your help!

Answer by fractalier(2101)

We set up a proportion...
3 / 26 = x / 39
Now cross-multiply and solve
26x = 117
x = 117/26 = 4 1/2 miles